OpenStudy (anonymous):

11) Express each of the following as indicated: a) 2 dm expressed in millimeters b) 2 h 10m expressed in seconds c) 16 g expressed in micrograms e) 0.675 mg expressed in grams f) 462 um expressed in centimeters g) 35/h expressed in meter per second

OpenStudy (anonymous):

Can somebody just help me with the explanation? I don't know what to do for e,f, &g.

OpenStudy (xishem):

We can just use conversion factors: e)\[0.675mg \times \frac{1g}{1000mg}=6.75 \times 10^{-4}g\] f)\[462 \mu m \times \frac{1m}{10^6 \mu m} \times \frac{100cm}{1m}=.462m\] I'm not sure what "h" is being referred to in b and g, though. Is there some context to the problem?

OpenStudy (shane_b):

For g since the answer is expressed in meters/second. However the numerator in the fraction doesn't show the units so that one is missing a key piece of info that's required to convert it.